# section[1.5em]1.5em1pc

### v2

@@ 1.   Introduction
Divided power structures are algebraic structures which are fairly ubiquitous when working over a field of positive characteristic. N. Jacobson introduced the concept of restricted Lie algebra to study modular Lie theory [Jacobson62]. A restricted Lie algebra is a Lie algebra equipped with an additional operation called the $p$-map, which satisfies specific relations. The archetypal example of a restricted Lie algebra is an associative algebra in prime characteristic equipped with the Lie bracket given by the commutator, and with the $p$-map given by the $p$-th power. Restricted Lie algebras appear notably in field theory [Jacobson37], linear algebraic groups [Borel91], and the cohomology of the Steenrod algebra [May66].
@@
The notion of algebra with divided powers was introduced by H. Cartan in [Cartan56exp7] to study certain bar constructions of commutative algebras, and further developed by N. Roby [Roby65]. A divided power algebra is a commutative algebra endowed with additional operations $\gamma_{n}$ which satisfy specific relations. An algebra $A$ over $\mathbb{Q}$ is naturally equipped with a structure of divided powers algebra such that $\gamma_{n}(x)=\frac{x^{n}}{n!}$ for all $x\in A$. Without any restriction on the ground ring, H. Cartan proves in [Cartan56exp7] that the homotopy of a simplicial commutative algebra comes naturally equipped with the structure of a divided power algebra. The notion of divided power algebra is essential in the theory of crystalline cohomology for schemes introduced by A. Grothendieck [Grothendieck68] and developed by P. Berthelot [Berthelot74]. The first author studied Quillen cohomology of divided power algebras [Dokas09, Dokas23].
@@
Operads are an algebraic device which represent types of algebras, such as associative algebras, commutative algebras, and Lie algebras. The theory of operads has proven to be a powerful tool in studying different categories of algebras in a unified way. For instance, in the seminal book [LodayV12], J.-L. Loday and B. Vallette make a detailed account of many known homological and homotopical theories for algebras over an operad. This includes a notion of Quillen (co)homology for algebras over an operad, studied in [Livernet98, Fresse98, Milles11, GoerssH00]. In the quadratic case, this is studied through a notion of Koszul duality due to V. Ginzburg and M. Kapranov [GinzburgK94] (see also [GetzlerJ94, Fresse04]). It also includes a notion of deformation theory due to M. Gerstenhaber [Gerstenhaber64] (see [Balavoine97, KontsevichS00, Keller05] for the operadic account), which led the way to the solution of the Deligne conjecture by D. Tamarkin [Tamarkin98].
The notion of divided power algebras over an operad was introduced by B. Fresse in [Fresse00]. While usual algebras over an operad $\mathcal{P}$ are algebras over a monad $S(\mathcal{P})$ built from $\mathcal{P}$ using coinvariant operations, divided power algebras over $\mathcal{P}$ are algebras over a monad $\Gamma(\mathcal{P})$ which is built using invariant operations. This recovers the classical notion of divided power algebras of Cartan, and Fresse shows that it also encompasses the notion of restricted Lie algebra of Jacobson. Furthermore, Fresse generalises Cartan’s result by showing that the homotopy of a simplicial algebra over $\mathcal{P}$ is always equipped with a structure of divided power algebra over $\mathcal{P}$. Divided power algebras were further studied by the third author, who obtained a convenient characterisation for these objects using monomial operations [Ikonicoff20].
The aim of this article is to study the Quillen cohomology of divided power algebras over any operad $\mathcal{P}$, using the techniques due to Quillen, Barr and Beck. This generalises work of the first author [Dokas04, Dokas09, Dokas23].
@@
I. General theory. In Sections 4 to 8, we develop the ingredients that go into Quillen cohomology for $\Gamma(\mathcal{P})$-algebras.
II. Examples. In Sections 9 and 10, we apply these constructions to the examples of classical divided power algebras, and of restricted Lie algebras, and recover known results.
III. Comparisons. In Sections 11 to 15, we investigate different comparison maps induced by adjunctions, which allow us to compare our cohomology for $\Gamma(\mathcal{P})$-algebras to more usual cohomology theories in certain categories of algebras.
@@
One side of the equivalence is obtained by building a new notion of semidirect product for $A$-modules.
In Section 6, we build a ring $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ which represents the operations defining the notion of an $A$-module. We obtain the following result:
Theorem (See 6.11).
@@
In Section 7, we identify Beck derivations in the category of $A$-modules, and we build an $A$-module $\Omega_{\Gamma(\mathcal{P})}(A)$ which represents these derivations.
@@
In Sections 9 and 10, we reconcile our general theory with the study of classical divided power algebras and of restricted Lie algebras by the first author in [Dokas04, Dokas09, Dokas15, Dokas23].
In the remaining sections of the article, we investigate the different comparison maps induced by certain adjunctions involving categories of divided power algebras. We focus on two types of such adjunctions. First, for any operad $\mathcal{P}$, there is an adjunction between the category of $\mathcal{P}$-algebras and the category of $\Gamma(\mathcal{P})$-algebras. Second, any morphism of operads $f\colon\mathcal{P}\to\mathcal{Q}$ induces an adjunction between the categories of $\Gamma(\mathcal{P})$- and $\Gamma(\mathcal{Q})$-algebras, similar to the extension/restriction of scalars adjunction for modules over a ring. In Section 11, we describe the comparison maps induced by both types of adjunctions in a general setting.
@@
Finally, in Section 15, we come back to the case were the base field is of characteristic 00, and we study the case of a good triple of operads $(\mathcal{C},\mathcal{A},\mathcal{P})$. In this case, the notions of $\mathcal{P}$-algebras and $\Gamma(\mathcal{P})$-algebras coincide. Good triples of operads generalise the mutual behaviours of the operads of commutative, of associative, and of Lie algebras. In particular, we see the operad $\mathcal{P}$ as a suboperad $\mathcal{A}$ of a certain type. In this setting, we can generalise the result of the previous section, and we obtain:
Theorem (See 15.4).
@@
Here again, $U(P)$ denotes a certain notion of universal enveloping $\mathcal{A}$-algebra, $\mathrm{HQ}_{\mathcal{A}\text{-}\mathrm{Alg}}^{*}$ is the Quillen cohomology for $\mathcal{A}$-algebras, and $\mathrm{HQ}_{\mathcal{P}\text{-}\mathrm{Alg}}^{*}$ is the Quillen cohomology for $\mathcal{P}$-algebras. If we consider the good triple $(\mathrm{Com},\operatorname{As},\operatorname{Lie})$, then 15.4 recovers the Cartan–Eilenberg result in characteristic zero.
While this last section focusses on the characteristic zero setting, there is evidence that the notion of good triple of operads can be generalised in positive characteristic, and we expect that our result still holds under the correct assumptions. Formulating these assumptions is—to our knowledge—an open problem for future consideration.
@@
All the operads considered in this article will be algebraic, symmetric operads. We assume the reader has a good familiarity with this notion of operad. For more details, we refer to [LodayV12].
@@
  • $\Gamma(\mathcal{P})\text{-}\mathrm{Alg}$ the category of $\Gamma(\mathcal{P})$-algebras in $\mathrm{Vect}_{\mathbb{F}}$, with forgetful functor $U^{\Gamma(\mathcal{P})}_{\mathcal{P}}\colon\Gamma(\mathcal{P})\text{-}\mathrm{Alg}\to\mathcal{P}\text{-}\mathrm{Alg}$ and its left adjoint $F^{\Gamma(\mathcal{P})}_{\mathcal{P}}\colon\mathcal{P}\text{-}\mathrm{Alg}\to\Gamma(\mathcal{P})\text{-}\mathrm{Alg}$.
@@ Acknowledgements
We thank Muriel Livernet for helpful discussions.
@@ 2.   Preliminaries on Quillen cohomology
Throughout this section, let $\mathbf{\mathcal{C}}$ be an algebraic category. More details about the kinds of categories we may consider are given in [Frankland15]*§2. For the purposes of this paper, the categories of $\mathcal{P}$-algebras and $\Gamma(\mathcal{P})$-algebras are examples of algebraic categories, in fact, the one-sorted kind: sets equipped with certain operations satisfying certain equations. Let us recall some terminology, also found in [Barr96] and [Frankland15]*§1.3.
@@
The abelianization over $X$ is the functor $Ab_{X}\colon\mathbf{\mathcal{C}}/X\to(\mathbf{\mathcal{C}}/X)_{\mathrm{ab}}$ which is left adjoint to the forgetful functor $U_{X}\colon(\mathbf{\mathcal{C}}/X)_{\mathrm{ab}}\to\mathbf{\mathcal{C}}/X$.
@@
The cotangent complex $\mathbf{L}_{X}$ of $X$ is the derived abelianization of $X$, i.e., the simplicial module over $X$ given by $\mathbf{L}_{X}\coloneqq Ab_{X}(C_{\bullet}\to X)$, where $C_{\bullet}\to X$ is a cofibrant replacement of $X$ in $s\mathbf{\mathcal{C}}$.
@@ 3.   Recollections on divided power algebras
In this section, we review the general notion of a divided power algebra over an operad $\mathcal{P}$, which are also called $\Gamma(\mathcal{P})$-algebras, due to Fresse [Fresse00]. Throughout this article, we use the characterisation of divided power algebras in terms of monomial operations due to the third author [Ikonicoff20]. We rely heavily on the notation introduced in [Ikonicoff20] for those operations, and operations on partitions of integers.
Recall that an operad $\mathcal{P}$ is a sequence $\{\mathcal{P}(n)\}_{n\in\mathbb{N}}$ such that for all $n$, $\mathcal{P}(n)$ is an $\mathbb{F}$-linear representation of the symmetric group $\Sigma_{n}$ on $n$ letters. An operad is called reduced when $\mathcal{P}(0)=0$, and Fresse showed the following:
@@
Here, $(\mathcal{P}(n)\otimes V^{\otimes n})^{\Sigma_{n}}$ stands for the module of invariant under the diagonal action of $\Sigma_{n}$, where the action of $\Sigma_{n}$ on $V^{\otimes n}$ is the usual permutation of factors.
@@
A divided power algebra over a reduced operad $\mathcal{P}$ is a vector space $A$ endowed with a family of operations $\beta_{x,{\underline{r}}}\colon A^{\times s}\to A$, given for all ${\underline{r}}=(r_{1},\dots,r_{s})$ such that $r_{1}+\dots+r_{s}=n$ and $x\in\mathcal{P}(n)^{\Sigma_{\underline{r}}}$, and which satisfy the relations:
@@
$\underline{Q}:=(\underbrace{p^{0},\dots,p^{0}}_{r_{1,0}},\dots,\underbrace{p^{k_{r_{1}}},\dots,p^{k_{r_{1}}}}_{r_{1,{k_{r_{1}}}}},\dots,\underbrace{p^{k_{r_{s}}},\dots,p^{k_{r_{s}}}}_{r_{s,{k_{r_{s}}}}}).$
@@
$\sum_{\sigma\in\Sigma_{\underline{R}}/\Sigma_{\underline{Q}}}\sigma x$ $=\left(\prod_{i=1}^{s}\prod_{j=1}^{k_{r_{i}}}\binom{r_{i,j}p^{j}}{p^{j},\dots,p_{j}}\right)x$
$=\left(\prod_{i=1}^{s}\prod_{j=1}^{k_{r_{i}}}r_{i,j}!\right)x.$
@@
$\beta_{x,\underline{R}}(\underbrace{a_{1},\dots,a_{1}}_{k_{r_{1}}+1},\dots,\underbrace{a_{s},\dots,a_{s}}_{k_{r_{s}}})=\left(\prod_{i=1}^{s}\prod_{j=1}^{k_{r_{i}}}\left(r_{i,j}!\right)^{-1}\right)\beta_{x,\underline{Q}}(\underbrace{a_{1},\dots,a_{1}}_{r_{1,0}+\dots+r_{1,k_{r_{1}}}},\dots,\underbrace{a_{s},\dots,a_{s}}_{r_{s,0}+\dots+r_{s,k_{r_{s}}}}).$
Here, the integers in $\underline{Q}$ are either 0s or powers of $p$. Using relation ($\beta$2) allows us to remove the 0s, and we obtain the result. ∎

### v3

@@ 1.   Introduction
Divided power structures are algebraic structures which are fairly ubiquitous when working over a field of positive characteristic. N. Jacobson introduced the concept of restricted Lie algebra to study modular Lie theory [Jacobson62]. A restricted Lie algebra is a Lie algebra equipped with an additional operation called the $p$-map, which satisfies specific relations. The archetypal example of a restricted Lie algebra is an associative algebra in prime characteristic equipped with the Lie bracket given by the commutator, and with the $p$-map given by the $p$-th power. Restricted Lie algebras appear notably in field theory [Jacobson37], linear algebraic groups [Borel91], and the cohomology of the Steenrod algebra [May66res].
@@
The notion of algebra with divided powers was introduced by H. Cartan in [Cartan56exp7] to study certain bar constructions of commutative algebras, and further developed by N. Roby [Roby65]. A divided power algebra is a commutative algebra endowed with additional operations $\gamma_{n}$ which satisfy specific relations. An algebra $A$ over $\mathbb{Q}$ is naturally equipped with a structure of divided powers algebra such that $\gamma_{n}(x)=\frac{x^{n}}{n!}$ for all $x\in A$. Without any restriction on the ground ring, H. Cartan proves in [Cartan56exp7] that the homotopy of a simplicial commutative algebra comes naturally equipped with the structure of a divided power algebra. Divided power algebras are essential in the theory of crystalline cohomology for schemes introduced by A. Grothendieck [Grothendieck68] and developed by P. Berthelot [Berthelot74]. The first author studied Quillen cohomology of divided power algebras [Dokas09, Dokas23].
@@
Operads are an algebraic device which represent types of algebras, such as associative algebras, commutative algebras, and Lie algebras. The theory of operads has proven to be a powerful tool in studying different categories of algebras in a unified way. For instance, in the seminal book [LodayV12], J.-L. Loday and B. Vallette make a detailed account of many known homological and homotopical theories for algebras over an operad. This includes Quillen (co)homology for algebras over an operad, studied in [Livernet98, Fresse98, Milles11, GoerssH00]. In the quadratic case, this is studied through a version of Koszul duality due to V. Ginzburg and M. Kapranov [GinzburgK94] (see also [GetzlerJ94, Fresse04]). It also includes a deformation theory due to M. Gerstenhaber [Gerstenhaber64]; see [Balavoine97, KontsevichS00, Keller05] for the operadic account. Homotopical resolutions of the Gerstenhaber operad led the way to D. Tamarkin’s solution of the Deligne conjecture [Tamarkin98]; see also [GerstenhaberV95hom, Hinich03].
The notion of divided power algebras over an operad was introduced by B. Fresse in [Fresse00]. While usual algebras over an operad $\mathcal{P}$ are algebras over a monad $S(\mathcal{P})$ built from $\mathcal{P}$ using coinvariant operations, divided power algebras over $\mathcal{P}$ are algebras over a monad $\Gamma(\mathcal{P})$ which is built using invariant operations. This recovers the classical divided power algebras of Cartan, and Fresse shows that it also encompasses the restricted Lie algebras of Jacobson. Furthermore, Fresse generalises Cartan’s result by showing that the homotopy of a simplicial algebra over $\mathcal{P}$ is always equipped with a structure of divided power algebra over $\mathcal{P}$. Divided power algebras were further studied by the third author, who obtained a convenient characterisation for these objects using monomial operations [Ikonicoff20].
The aim of this article is to study the Quillen cohomology of divided power algebras over any operad $\mathcal{P}$, using techniques due to Quillen, Barr and Beck. This generalises work of the first author [Dokas04, Dokas09, Dokas23].
@@
Part I. General theory. In Sections 4 to 8, we develop the ingredients that go into Quillen cohomology for $\Gamma(\mathcal{P})$-algebras.
@@
One side of the equivalence is obtained by building a semidirect product for $A$-modules.
In Section 6, we build a ring $\mathbb{U}_{\Gamma(\mathcal{P})}(A)$ which represents the operations that define an $A$-module. We obtain the following result:
Theorem (See Theorem 6.11).
@@
In Section 7, we identify Beck derivations in the category of $A$-modules, and we construct an $A$-module $\Omega_{\Gamma(\mathcal{P})}(A)$ which represents these derivations.
@@
Part II. Examples. In Sections 9 and 10, we apply our general theory to the examples of classical divided power algebras and of restricted Lie algebras. We recover some results obtained by the first author in [Dokas04, Dokas09, Dokas15, Dokas23].
Part III. Comparisons. In Sections 11 to 15, we investigate different comparison maps induced by adjunctions, which allow us to compare our cohomology for $\Gamma(\mathcal{P})$-algebras to more usual cohomology theories in certain categories of algebras.
We focus on two types of adjunctions involving divided power algebras. First, for any operad $\mathcal{P}$, there is an adjunction between the category of $\mathcal{P}$-algebras and the category of $\Gamma(\mathcal{P})$-algebras. Second, any morphism of operads $f\colon\mathcal{P}\to\mathcal{Q}$ induces an adjunction between the categories of $\Gamma(\mathcal{P})$- and $\Gamma(\mathcal{Q})$-algebras, similar to the extension/restriction of scalars adjunction for modules over a ring. In Section 11, we describe the comparison maps induced by both types of adjunctions in a general setting.
@@
Finally, in Section 15, we come back to the case where the base field is of characteristic 00, and we study the case of a good triple of operads $(\mathcal{C},\mathcal{A},\mathcal{P})$. In this case, the notions of $\mathcal{P}$-algebras and $\Gamma(\mathcal{P})$-algebras coincide. Good triples of operads generalise the mutual behaviours of the operads of commutative, of associative, and of Lie algebras. In particular, we see the operad $\mathcal{P}$ as a suboperad $\mathcal{A}$ of a certain type. In this setting, we can generalise the result of the previous section, and we obtain:
Theorem (See Theorem 15.4).
@@
Here again, $U(P)$ denotes a certain notion of universal enveloping $\mathcal{A}$-algebra, $\mathrm{HQ}_{\mathcal{A}\text{-}\mathrm{Alg}}^{*}$ is the Quillen cohomology for $\mathcal{A}$-algebras, and $\mathrm{HQ}_{\mathcal{P}\text{-}\mathrm{Alg}}^{*}$ is the Quillen cohomology for $\mathcal{P}$-algebras. If we consider the good triple $(\mathrm{Com},\operatorname{As},\operatorname{Lie})$, then Theorem 15.4 recovers the Cartan–Eilenberg result in characteristic zero.
While this last section focusses on the characteristic zero setting, there is evidence that good triples of operads can be generalised in positive characteristic, and we expect that our result still holds under the correct assumptions. Formulating these assumptions is—to our knowledge—an open problem for future consideration.
@@
All the operads considered in this article will be algebraic, symmetric operads. We assume the reader has a good familiarity with this type of operad. For more details, we refer to [LodayV12].
@@
  • $\Gamma(\mathcal{P})\text{-}\mathrm{Alg}$ the category of $\Gamma(\mathcal{P})$-algebras in $\mathrm{Vect}_{\mathbb{F}}$, with forgetful functor $U^{\Gamma(\mathcal{P})}_{\mathcal{P}}\colon\Gamma(\mathcal{P})\text{-}\mathrm{Alg}\to\mathcal{P}\text{-}\mathrm{Alg}$ and its left adjoint functor, $F^{\Gamma(\mathcal{P})}_{\mathcal{P}}\colon\mathcal{P}\text{-}\mathrm{Alg}\to\Gamma(\mathcal{P})\text{-}\mathrm{Alg}$.
@@ Acknowledgements
We thank Muriel Livernet for helpful discussions. We also thank the anonymous referee for their useful suggestions.
@@ 2.   Preliminaries on Quillen cohomology
Throughout this section, let $\mathbf{\mathcal{C}}$ be an algebraic category, i.e., a category of models for a finite product sketch (or equivalently, for a many-sorted Lawvere theory). More details about algebraic categories are given in [Frankland15]*§2. For the purposes of this paper, the categories of $\mathcal{P}$-algebras and $\Gamma(\mathcal{P})$-algebras are examples of algebraic categories, in fact, the one-sorted kind: sets equipped with certain operations satisfying certain equations. Let us recall some terminology, also found in [Barr96] and [Frankland15]*§1.3.
@@
The abelianization over $X$ is the left adjoint functor $Ab_{X}\colon\mathbf{\mathcal{C}}/X\to(\mathbf{\mathcal{C}}/X)_{\mathrm{ab}}$ to the forgetful functor $U_{X}\colon(\mathbf{\mathcal{C}}/X)_{\mathrm{ab}}\to\mathbf{\mathcal{C}}/X$.
@@
Note that abelianizations and pushforwards exist when $\mathbf{\mathcal{C}}$ is an algebraic category [Frankland15]*Proposition 3.26.
@@
The cotangent complex $\mathbf{L}_{X}$ of $X$ is the derived abelianization of $X$, i.e., the simplicial module over $X$ given by $\mathbf{L}_{X}\coloneqq Ab_{X}(C_{\bullet}\to X)$, where $C_{\bullet}\xrightarrow{\sim}X$ is a cofibrant replacement of $X$ in $s\mathbf{\mathcal{C}}$, the category of simplicial objects in $\mathbf{\mathcal{C}}$.
Here $s\mathbf{\mathcal{C}}$ is endowed with the standard Quillen model structure [Quillen67]*§II.4 [GoerssJ09]*§II.5.
@@ 3.   Recollections on divided power algebras
In this section, we review the general notion of a divided power algebra over an operad $\mathcal{P}$, also called a $\Gamma(\mathcal{P})$-algebra, due to Fresse [Fresse00]. Throughout this article, we use the characterisation of divided power algebras in terms of monomial operations due to the third author [Ikonicoff20]. We rely heavily on the notation introduced in [Ikonicoff20] for those operations, and operations on partitions of integers.
Recall that an operad $\mathcal{P}$ is (in particular) a sequence of $\mathbb{F}$-vector spaces $\{\mathcal{P}(n)\}_{n\in\mathbb{N}}$ such that $\mathcal{P}(n)$ has a right action of $\Sigma_{n}$, the symmetric group on $n$ letters. An operad is called reduced when $\mathcal{P}(0)=0$, and Fresse showed the following:
@@
Here, $(\mathcal{P}(n)\otimes V^{\otimes n})^{\Sigma_{n}}$ stands for the module of invariants under the diagonal action of $\Sigma_{n}$, where the action of $\Sigma_{n}$ on $V^{\otimes n}$ is the usual permutation of factors.
@@
A divided power algebra over a reduced operad $\mathcal{P}$ is a vector space $A$ endowed with a family of operations $\beta_{x,{\underline{r}}}\colon A^{\times s}\to A$, given for all ${\underline{r}}=(r_{1},\dots,r_{s})$ such that $r_{1}+\dots+r_{s}=n$ and $x\in\mathcal{P}(n)^{\Sigma_{\underline{r}}}$, which satisfy the relations:
@@
$\underline{Q}\coloneqq(\underbrace{p^{0},\dots,p^{0}}_{r_{1,0}},\dots,\underbrace{p^{k_{r_{1}}},\dots,p^{k_{r_{1}}}}_{r_{1,{k_{r_{1}}}}},\dots,\underbrace{p^{k_{r_{s}}},\dots,p^{k_{r_{s}}}}_{r_{s,{k_{r_{s}}}}}).$
@@
$\sum_{\sigma\in\Sigma_{\underline{R}}/\Sigma_{\underline{Q}}}\sigma x$ $=\left(\prod_{i=1}^{s}\prod_{j=1}^{k_{r_{i}}}\binom{r_{i,j}p^{j}}{p^{j},\dots,p_{j}}\right)x=\left(\prod_{i=1}^{s}\prod_{j=1}^{k_{r_{i}}}r_{i,j}!\right)x.$
@@
$\beta_{x,\underline{R}}(\underbrace{a_{1},\dots,a_{1}}_{k_{r_{1}}+1},\dots,\underbrace{a_{s},\dots,a_{s}}_{k_{r_{s}}})=\\ \left(\prod_{i=1}^{s}\prod_{j=1}^{k_{r_{i}}}\left(r_{i,j}!\right)^{-1}\right)\beta_{x,\underline{Q}}(\underbrace{a_{1},\dots,a_{1}}_{r_{1,0}+\dots+r_{1,k_{r_{1}}}},\dots,\underbrace{a_{s},\dots,a_{s}}_{r_{s,0}+\dots+r_{s,k_{r_{s}}}}).$
Here, the integers in $\underline{Q}$ are either 00 or powers of $p$. Using relation ($\beta$2) allows us to remove the zeros, and we obtain the result. ∎